QUANTIZATION OF PRESYMPLECTIC MANIFOLDS AND CIRCLE ACTIONS

ANA CANAS DA SILVA YAEL KARSHON AND SUSAN TOLMAN

Abstract. We prove several versions of “quantization commutes with reduction” for circle actions on manifolds that are not symplectic. Instead, these manifolds possess a weaker structure, such as a spin^{c} structure. Our theorems work whenever the quantization data and the reduction data are compatible; this condition always holds if we start from a presymplectic (in particular, symplectic) manifold.